FIFO, LIFO, Weighted Average

Date

# units purchased

Cost / Unit

# of units sold

Selling price/ unit

January 1

Beg. Inventory

200 Units

$10 / unit

January 10

Purchase

200 Units

$12 / unit

January 15

Purchase

200 Units

$14 / unit

January 31

Sold

250 units

$18 / unit

Prepare JE to record January 31 sale assuming:

  1. FIFO
  2. LIFO
  3. Weighted Average

Which method best matches most recent costs to Net Sales?

Which method produces an Inventory value that is closest to current replacement costs?
Answer #1

1.) Journal entry for sale of goods

a) FIFO

Cost of goods sold = (200*10) +(50*12) = 2,000+600 = 2,600

Sale Price = 250*18 = 4,500

Entry

Dr Cash 4,500

Dr Cost of Goods Sold 2,600

Cr Sales 4,500

Cr Inventory 2,600

b) LIFO

Cost of goods sold = (200*14) + (50*12) = 2,800 + 600 = 3,400

Sale price = 250*18 = 4,500

Entry

Dr Cash 4,500

Dr Cost of Goods Sold 3,400

Cr Sales 4,500

Cr Inventory 3,400

c) Weighted Average Price Method

Cost of goods sold per unit = {(200*10)+(200*12)+(200*14)}/(200+200+200)

= (2,000+2,400+2,800) / 600

= 7,200/600 = $12

Cost of goods sold = (250*12) = 3,000

Sale price = 250*18 = 4,500

Entry

Dr Cash 4,500

Dr Cost of Goods Sold 3,000

Cr Sales 4,500

Cr Inventory 3,000

Note : It is assumed that the sales are made for cash. If it is made for credit, then corresponding "Debtor" account will be debited instead of Cash

2.) Which method best matches most recent costs to Net Sales?

Ans: LIFO. As this method considers the cost of goods sold to be the newest purchased inventory cost.

3.) Which method produces an Inventory value that is closest to current replacement costs?

Ans : FIFO. As the cost of goods sold are computed as the cost of oldest inventory purchased, the closing inventory will contain the cost of newest inventory purchased which is at current replacement cost.